# Phonon-mediated spin-polarized superconductivity in altermagnets

### v1

@@ I Introduction
More than a century after its discovery [1, 2], superconductivity continues to be one of the most intensely studied topics in condensed matter physics, with deep connections to the most fundamental descriptions of matter [3, 4, 5, 6]. This macroscopic quantum phenomenon is characterized by zero electrical resistance and Higgs condensation rendering the photon massive [7, 3, 5] below some critical temperature $T_{c}$. Superconductors arising out of good metals with small correlation effects (conventional low-$T_{c}$ superconductors with a nodeless $s$-wave superconducting gap), are well described by the Bardeen-Cooper-Schrieffer (BCS) theory [8]. Within BCS theory the phenomenon arises from an instability of the Fermi surface (FS) due to effective attraction between electrons. Originally, exchange of phonons mediated the attraction. The discovery of unconventional superconductivity with nodal gaps in strongly correlated fermion systems such as heavy fermions [9, 10] and high-$T_{c}$ superconductivity [11, 12, 13, 14, 15] highlighted that other bosons could also be responsible for the pairing. In unconventional superconductors [16], the pairing mechanism often involves complex interactions, such as spin fluctuations, electronic correlations, or orbital effects, leading to non-trivial symmetry and momentum-dependent superconducting gaps. In the high-$T_{c}$ cuprates, nodes in the gap on the FS have been established through phase-sensitive measurements [17], establishing that the gap is spin-singlet with $d_{x^{2}-y^{2}}$-wave symmetry. Furthermore, low-$T_{c}$, possibly $p$-wave, spin-triplet pairing in itinerant ferromagnets has been predicted and observed [18, 19, 20, 21, 22]. Finally, magnon-mediated unconventional superconductivity in heterostructures of magnetic insulators and various gapless fermionic systems has been considered extensively [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
@@
Recently, altermagnets have emerged as a new class of magnetic materials that defy the traditional dichotomy of collinear ferromagnetism and antiferromagnetism [38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Altermagnets exhibit time-reversal symmetry breaking and spin-split band structures, yet maintain compensated magnetic ordering without net magnetization. They combine properties of ferromagnets and antiferromagnets, opening new avenues for superconductivity research [48, 49, 50, 51, 52, 53, 39, 54]. Importantly, the spin-split electronic bands in altermagnets exhibit symmetries that can support exotic pairing states in the presence of suitable interactions. However, spin-singlet or mixed-spin spin-triplet superconductivity requires finite-momentum Cooper pairs or odd-frequency superconductivity [55, 51, 50]. Even-frequency zero-momentum Cooper-pairing requires fully spin-polarized Cooper pairs.
@@
Previously, unconventional magnon-mediated superconductivity has been investigated in a minimal model of altermagnets [39] on a Lieb-lattice [56]. This particular model exhibits antiferromagnetism along with $d_{x^{2}-y^{2}}$-wave symmetric spin splitting of the electron bands. The existence of altermagnetism in Lieb lattices has also been confirmed using Hubbard models [57, 58]. In addition, several proposed quasi-two-dimensional (2D) altermagnets display the Lieb lattice [58, 46, 47, 59, 60]. Magnon-mediated superconductivity with zero-momentum Cooper-pairs in altermagnets, requires invoking double-magnon processes to obtain superconductivity [39]. In a strong-coupling superconductivity scenario, many-body effects on this superconductivity has been investigated [54] using Eliashberg theory of strong-coupling superconductivity [61, 62]. In this case, strong-coupling renormalizations leads to substantial suppression of the critical temperature $T_{c}$ [54]. Were the electron-electron interaction to be mediated by a spinless boson, the necessity of a double-boson mediated interaction would be obviated, and the strong-coupling reduction of $T_{c}$ possibly avoided.
@@
In this work, we predict the emergence of unconventional fully spin-polarized superconductivity in altermagnets mediated by electron-phonon interactions. The examples mentioned above of phonon-mediated conventional and unconventional spin-singlet superconductivity are not an option in the spin-split band structure of altermagnets, at least not for zero-momentum Cooper-pairs. Our result originates from the interplay between spin-symmetries of the electron bands in altermagnets, and the momentum dependence of the electron-phonon interaction. This provides a new pathway for realizing unconventional superconductivity via a conventional mechanism, and with substantial $T_{c}$.
@@ II Phonon model
$\sum_{\beta}D^{\alpha\beta}(\textbf{q})\bm{\xi}^{\beta}(\textbf{q})=M_{\alpha}\omega^{2}(\textbf{q})\bm{\xi}^{\alpha}(\textbf{q}),$ (1)
where $\alpha,\beta\in\{1,2,3\}$ indicate the different atoms in the basis of the Lieb lattice in Fig. 1 (a), referring to the spin-up, non-magnetic, and spin-down sublattices respectively, $D$ is the dynamical matrix, which is the Fourier transformation of the force constant matrix (FCM), $\bm{\xi}(\textbf{q})$ is the Fourier transformation of the displacement vector, q is the momentum of the phonon, $M_{\alpha}$ is the mass of atom $\alpha$, and $\omega(\textbf{q})$ is the phonon frequency (see Supplemental Material [70] for details). To lowest order in displacements, only in-plane phonons contribute to the electron-phonon coupling in a 2D system, hence, we exclude the out-of-plane modes. The FCM shares the eight symmetries of the Lieb lattice [39, 70], which will put restrictions on the elements of the FCM. However, there are still $11$ free parameters, nevertheless, by imposing additional constraints, we have only one free parameter, which we set to give reasonable phonon bands shown in Fig. 1 (b) (see Supplemental Material for details [70]).
Figure 1: (a) The Lieb lattice contains three atoms in the basis with different localized spins, namely, spin-up (red), non-magnetic (yellow), and spin-down (blue). (b) The in-plane phonon bands for the Lieb lattice on a path between the symmetry points of the Brillouin zone, see Supplemental Material [70] for the parameter values used.
@@ III Electron model
where $c_{i\sigma}^{({\dagger})}$ annihilates (creates) an electron at site $i$ with spin $\sigma$, the hopping amplitude between NNs, $\langle i,j\rangle$, and NNNs, $\langle\langle i,j\rangle\rangle$, are given by $t$ and $t_{2}$ respectively, and $J_{\mathrm{sd}}$ is the interaction strength between the itinerant electrons and the on-site spins, $\textbf{S}_{i}$, with $\bm{\sigma}$ being the vector consisting of Pauli matrices. The spins $\textbf{S}_{i}$ have a length $S$ and the on-site energy is given by the chemical potential, $\mu$.
The electronic energy spectrum obtained from this Hamiltonian is showcased in Fig. 2 (a), which exhibits the characteristic $d_{x^{2}-y^{2}}$-wave spin-split electron bands of an altermagnet. To obtain this spectrum, we diagonalize the Hamiltonian by turning to momentum space as $H_{\mathrm{el}}=\sum_{n\textbf{k}\sigma}\varepsilon_{n\textbf{k}\sigma}d^{\dagger}_{n\textbf{k}\sigma}d_{n\textbf{k}\sigma}$, where $\varepsilon_{n\textbf{k}\sigma}$ is the energy of band $n$ for excitations with momentum k and spin $\sigma$, while $d^{\left({\dagger}\right)}_{n\textbf{k}\sigma}$ are the annihilation (creation) band operators. These operators are related to the electron operators as $d_{n\textbf{k}\sigma}=\sum_{\alpha}q_{\textbf{k}\sigma\alpha n}^{*}c_{\textbf{k}\sigma\alpha}$, where $\alpha$ indicates the different atoms in the basis, and the coefficient $q_{\textbf{k}\sigma\alpha n}$ is the component $\alpha$ in the eigenvector of the Hamiltonian corresponding to band $n$.
Due to the $d_{x^{2}-y^{2}}$-wave symmetry of the spin splitting, spin-up and spin-down bands are interchangeable by a rotation of $90^{\circ}$ in momentum space. Hence, all properties of the two spins are connected by this rotation. To showcase the role of NNN hopping, we present the results with a NNN hopping of $t_{2}\approx 0.21t$ for spin-up electrons and the results with $t_{2}=0$ for spin-down electrons throughout the letter. The electronic band structure for $t_{2}=0$ can be found in Ref. [39].
Figure 2: (a) The electronic band structure for the two spins along a path between the symmetry points in the Brillouin zone. The values used are $t=2.4$ eV, $t_{2}\approx 0.21t$ (for both spins), $J_{\mathrm{sd}}S=0.7$ eV, and $\mu=0$. (b) The Fermi surfaces for spin-up (red) and spin-down (blue), the colors indicate the inverse derivative of the bands perpendicular at the Fermi surface, $\left|d\varepsilon/dk_{\perp}\right|^{-1}$. The values used are $t=2.4$ eV, $J_{\mathrm{sd}}S=0.7$ eV, $\mu=0.1t$, $t_{2}\approx 0.21t$ for spin-up, and $t_{2}=0$ for spin-down. This particular system exhibits altermagnetism with $d_{x^{2}-y^{2}}$-wave spin splitting.
@@ IV Electron-phonon coupling
where $\textbf{r}_{i\alpha j\beta}$ is the difference between the instantaneous position vectors for atom $i\alpha$ and $j\beta$, $\bm{\delta}$ are vectors connecting the atoms in equilibrium, $\textbf{u}_{ij\alpha\beta}$ is the difference in displacement from equilibrium of atom $i\alpha$ and $j\beta$, $\bm{\nabla}_{r}$ is the gradient with respect to $\textbf{r}_{i\alpha j\beta}$, and $|_{\textbf{r}_{ij\alpha\beta}=\bm{\delta}}$ means to evaluate the atom positions in equilibrium (see Supplemental Material [70]). The first term of (3) is the equilibrium hopping used in (2), while the second term is the source of the electron-phonon interaction. We quantize the displacements as [66, 67]
$\textbf{u}_{i\alpha}=\sum_{\textbf{q}\lambda}\sqrt{\frac{\hbar}{2M_{\alpha}N\omega_{\textbf{q}\lambda}}}\bm{\hat{e}}_{\textbf{q}\alpha}^{\lambda}(a_{-\textbf{q},\lambda}^{\dagger}+a_{\textbf{q}\lambda})e^{-i\textbf{q}\cdot\textbf{R}_{i\alpha}},$ (4)
where $a_{\textbf{q}\lambda}^{({\dagger})}$ annihilates (creates) a phonon with momentum q in mode $\lambda$, $M_{\alpha}$ is the mass of atom $\alpha$ in the basis, $N$ is the number of unit cells, $\omega_{\textbf{q}\lambda}$ is the frequency of a phonon in mode $\lambda$ with momenta q, $\bm{\hat{e}}_{\textbf{q}\alpha}^{\lambda}$ is the normalized eigenvector of the dynamical matrix for a phonon in mode $\lambda$ with momentum q corresponding to the displacement of atom $\alpha$, and $\textbf{R}_{i\alpha}$ is the equilibrium position of atom $\alpha$ in unit cell $i$.
@@
×$\times\sum_{\bm{\delta}_{j}}e^{i\textbf{k}\cdot\bm{\delta}_{j}}\left(\frac{\delta_{jx}}{\sigma_{x}^{2}},\frac{\delta_{jy}}{\sigma_{y}^{2}}\right)\cdot\left[\frac{\bm{\hat{e}}_{\textbf{q}\beta}^{\lambda}}{\sqrt{M_{\beta}}}e^{i\textbf{q}\cdot\bm{\delta}_{j}}-\frac{\bm{\hat{e}}_{\textbf{q}\alpha}^{\lambda}}{\sqrt{M_{\alpha}}}\right],$
where $j\in\{1,2\}$ indicates whether it is the NN or the NNN coupling. Note that we have omitted the band indices, as we consider the scattering to be confined to the FS.
@@ V Superconductivity
Next, we transform the electron-phonon interaction into an effective electron-electron interaction by applying a Schrieffer-Wolff transformation [72, 73] (see Supplemental Material [70])
$H_{\mathrm{eff}}=\frac{1}{2}\sum_{\textbf{k}\textbf{k}^{\prime}\sigma}\bar{V}_{\textbf{k}^{\prime}\textbf{k}\sigma}d^{\dagger}_{\textbf{k}^{\prime}\sigma}d^{\dagger}_{-\textbf{k}^{\prime},\sigma}d_{-\textbf{k},\sigma}d_{\textbf{k}\sigma},$ (7)
where we have followed Refs. [34, 54], and symmetrized the effective interaction as $2\bar{V}_{\textbf{k}^{\prime}\textbf{k}\sigma}\equiv V_{\textbf{k}^{\prime}\textbf{k}\sigma}+V_{-\textbf{k}^{\prime},-\textbf{k}\sigma}-V_{-\textbf{k}^{\prime},\textbf{k}\sigma}-V_{\textbf{k}^{\prime},-\textbf{k}\sigma}$, with the effective electron-electron interaction on the FS of spin $\sigma$, $V_{\textbf{k}^{\prime}\textbf{k}\sigma}$ given by
@@
The symmetrized interaction is shown in Fig. 3 (a) and depicts a $p$-wave symmetry. The electron band structure in an altermagnet rules out spin-singlet zero-momentum pairing. Hence, we look for spin-triplet pairing which necessitates odd-parity gap functions in the case of even-frequency pairing [55], and hence an odd-parity symmetrized coupling [16, 34]. The linearized self-consistent equation for the spin-polarized superconducting gap $\Delta_{\textbf{k}\sigma}$, is given by $\Delta_{\textbf{k}\sigma}=-\sum_{\textbf{k}^{\prime}}\bar{V}_{\textbf{k}\textbf{k}^{\prime}\sigma}\frac{\Delta_{\textbf{k}^{\prime}\sigma}}{2\left|\varepsilon_{\textbf{k}^{\prime}\sigma}\right|}\tanh(\beta\left|\varepsilon_{\textbf{k}^{\prime}\sigma}\right|/2)$. If we momentarily consider a jellium model for phonons [67], the strength of the electron-phonon coupling depends only on the momentum transfer. Still considering spin-polarized superconductivity, the symmetrized coupling would be $\bar{V}_{\textbf{k}\textbf{k}^{\prime}\sigma}=-|g_{\textbf{k}-\textbf{k}^{\prime}}|^{2}/\hbar\omega_{\textbf{k}-\textbf{k}^{\prime}}+|g_{\textbf{k}+\textbf{k}^{\prime}}|^{2}/\hbar\omega_{\textbf{k}+\textbf{k}^{\prime}}$ on the FS for spin $\sigma$.
Let us first consider a simple model for acoustic phonons, $g_{\textbf{q}}=g\sqrt{|\textbf{q}|}$ and $\omega_{\textbf{q}}=u|\textbf{q}|$. Then, $|g_{\textbf{q}}|^{2}/\hbar\omega_{\textbf{q}}=g^{2}/\hbar u$ is momentum independent, and so $\bar{V}_{\textbf{k}\textbf{k}^{\prime}\sigma}=0$. There is no phonon-mediated spin-polarized Cooper pairing from acoustic phonons within such a simple continuum model. Next, consider a simple model of Einstein phonons, $g_{\textbf{q}}=g|\textbf{q}|$ and $\omega_{\textbf{q}}=\omega_{E}$. Then, $\bar{V}_{\textbf{k}\textbf{k}^{\prime}\sigma}=g^{2}(|\textbf{k}+\textbf{k}^{\prime}|^{2}-|\textbf{k}-\textbf{k}^{\prime}|^{2})/\hbar\omega_{E}.$ Imagining an approximately circular FS around the $\bm{\Gamma}$ point, the symmetrized interaction closely resembles $\bar{V}_{\textbf{k}\textbf{k}^{\prime}\sigma}=V\cos\phi_{\textbf{k}\textbf{k}^{\prime}}$, with $\phi_{\textbf{k}\textbf{k}^{\prime}}$ being the angle between k and $\textbf{k}^{\prime}$, and $V>0$ some positive energy. When solving the eigenvalue problem in (9) for a positive eigenvalue $\lambda$, it is advantageous for the interaction to be attractive when $\textbf{k}^{\prime}$ is close to k, due to the negative sign on the RHS. A coupling of the form $\bar{V}_{\textbf{k}\textbf{k}^{\prime}\sigma}=V\cos\phi_{\textbf{k}\textbf{k}^{\prime}}$ with $V>0$ prevents non-trivial gap solutions, since the strongest coupling is always at $\textbf{k}=\pm\textbf{k}^{\prime}$, with a disadvantageous sign. This illustrates the key difference between phonon- and magnon-mediated interactions, since for magnons typically $|g_{\textbf{k}-\textbf{k}^{\prime}}|>|g_{\textbf{k}+\textbf{k}^{\prime}}|$ when $\textbf{k}\approx\textbf{k}^{\prime}$ due to the strength of the interaction increasing for long-wavelength magnons [27]. In contrast, for long-wavelength optical phonons, the coupling tends to vanish. Hence, in simple continuum models, optical phonons are also unlikely to yield spin-polarized superconductivity.
The electron-phonon coupling in (6) depends on both the incoming and outgoing momentum. The simple forms of $g_{\textbf{q}}$ and $\omega_{\textbf{q}}$ of the jellium model are replaced by complex phase factors giving a richer momentum structure on a lattice. Thus, both acoustic and optical phonons contribute to the symmetrized pairing, giving a non-zero symmetrized interaction in Fig. 3 (a). While the signs are still disadvantageous at $\textbf{k}=\pm\textbf{k}^{\prime}$, the momentum dependence is sufficiently anisotropic to allow solutions of the gap equation with an appreciable $T_{c}$, see below.
Using the phonon-mediated interaction of (7) for the altermagnet, we consider weak-coupling superconductivity and apply the linearized BCS equation to solve for the gap and critical temperature. Since electrons close to the FS dominate the superconducting pairing, we utilize the FS average of the BCS equation [39, 54, 74, 75]

### v2

@@ I Introduction
More than a century after its discovery [1, 2], superconductivity continues to be one of the most intensely studied topics in condensed matter physics, with deep connections to the most fundamental descriptions of matter [3, 4, 5, 6]. This macroscopic quantum phenomenon is characterized by zero electrical resistance and Higgs condensation, rendering the photon massive [7, 3, 5] below some critical temperature $T_{c}$. Superconductors arising out of good metals with small correlation effects (conventional low-$T_{c}$ superconductors with a nodeless $s$-wave superconducting gap), are well described by the Bardeen-Cooper-Schrieffer (BCS) theory [8]. Within BCS theory the phenomenon arises from an instability of the Fermi surface (FS) due to effective attraction between electrons. Originally, exchange of phonons mediated the attraction. The discovery of unconventional superconductivity with nodal gaps in strongly correlated fermion systems such as heavy fermions [9, 10] and high-$T_{c}$ superconductivity [11, 12, 13, 14, 15] highlighted that other bosons could also be responsible for the pairing. In unconventional superconductors [16], the pairing mechanism often involves complex interactions, such as spin fluctuations, electronic correlations, or orbital effects, leading to non-trivial symmetry and momentum-dependent superconducting gaps. In the high-$T_{c}$ cuprates, nodes in the gap on the FS have been established through phase-sensitive measurements [17], establishing that the gap is spin-singlet with $d_{x^{2}-y^{2}}$-wave symmetry. Furthermore, low-$T_{c}$, possibly $p$-wave, spin-triplet pairing in itinerant ferromagnets has been predicted and observed [18, 19, 20, 21, 22]. Finally, magnon-mediated unconventional superconductivity in heterostructures of magnetic insulators and various gapless fermionic systems has been considered extensively [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
@@
Recently, altermagnets have emerged as a new class of magnetic materials that defy the traditional dichotomy of collinear ferromagnetism and antiferromagnetism [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. Altermagnets exhibit time-reversal symmetry breaking and spin-split band structures, yet maintain compensated magnetic ordering without net magnetization. They combine properties of ferromagnets and antiferromagnets, opening new avenues for superconductivity research [51, 52, 53, 54, 55, 56, 57, 39, 58]. Importantly, the spin-split electronic bands in altermagnets exhibit symmetries that can support exotic pairing states [59] in the presence of suitable interactions. However, spin-singlet or mixed-spin spin-triplet superconductivity requires finite-momentum Cooper pairs or odd-frequency superconductivity [60, 55, 54]. Even-frequency zero-momentum Cooper-pairing requires fully spin-polarized Cooper pairs.
@@
Previously, unconventional magnon-mediated superconductivity has been investigated in a minimal model of altermagnets [39, 58] on a Lieb lattice [61, 62, 63, 64, 65]. This particular model exhibits antiferromagnetism along with $d_{x^{2}-y^{2}}$-wave symmetric spin splitting of the electron bands. The existence of altermagnetism in Lieb lattices has also been confirmed using Hubbard models [66, 67]. While realizing a 2D electronic Lieb lattice has been a longstanding challenge [64, 65], several proposed quasi-two-dimensional (2D) altermagnets display the Lieb lattice in nearly decoupled layers [67, 46, 47, 68, 48]. Magnon-mediated superconductivity with zero-momentum Cooper-pairs in altermagnets, requires invoking double-magnon processes to obtain superconductivity [39, 58]. In a strong-coupling superconductivity scenario, many-body effects on this superconductivity has been investigated [58] using Eliashberg theory of strong-coupling superconductivity [69, 70]. In this case, strong-coupling renormalizations leads to substantial suppression of the critical temperature $T_{c}$ [58]. Were the electron-electron interaction to be mediated by a spinless boson, the necessity of a double-boson mediated interaction would be obviated, and the strong-coupling reduction of $T_{c}$ possibly avoided.
@@
In this work, we predict the emergence of unconventional fully spin-polarized superconductivity in altermagnets mediated by electron-phonon interactions. The examples mentioned above of phonon-mediated conventional and unconventional spin-singlet superconductivity are not an option in the spin-split band structure of altermagnets, at least not for zero-momentum Cooper-pairs. Spin-polarized supercurrents are attractive for the field of superconducting spintronics [74], usually realized in complicated heterostructures relying on proximity induced superconductivity. We propose that spin polarized phonon-mediated superconductivity can arise intrinsically in altermagnets. Our result originates from the interplay between spin-symmetries of the electron bands in altermagnets, and the momentum dependence of the electron-phonon interaction. This provides a new pathway for realizing unconventional superconductivity via a conventional mechanism, and with a substantial $T_{c}$.
@@ II Phonon model
$\sum_{\beta}D^{\alpha\beta}_{\textbf{q}}\bm{\hat{e}}_{\textbf{q}\beta}^{\lambda}=M_{\alpha}\omega_{\textbf{q}\lambda}^{2}\bm{\hat{e}}_{\textbf{q}\alpha}^{\lambda},$ (1)
where $\alpha,\beta\in\{1,2,3\}$ label the three atoms in the Lieb lattice basis illustrated in Fig. 1 (a), referring to the spin-up, non-magnetic, and spin-down sublattices respectively. $D$ is the dynamical matrix, which is the Fourier transformation of the force constant matrix (FCM). The phonon modes are indicated by $\lambda$, $\bm{\hat{e}}_{\textbf{q}}$ is the normalized vector of the Fourier transformed displacement vector, q is the momentum of the phonon, $M_{\alpha}$ is the mass of atom $\alpha$, and $\omega_{\textbf{q}\lambda}$ is the phonon frequency. The symmetries of the Lieb lattice, given in App. A, constrain the elements of the FCM [39], leaving $11$ free parameters. As shown in App. B, by applying additional constraints, we are able to relate these parameters and are only left with one free parameter, $\eta$. This parameter will only affect the phonon bandwidth, which we have chosen to be around $30$ meV. As shown in App. C, in our first-order electron-phonon coupling in a 2D system, only the in-plane (IP) phonon modes will couple to the electrons, as such, we discard the out-of-plane (OOP) modes. The IP bands are shown in Fig. 1 (b).
Figure 1: (a) The Lieb lattice, where the black square indicates the chosen unit cell, which contains three atoms with different localized spins, namely, spin-up (red), non-magnetic (yellow), and spin-down (blue). (b) The IP phonon bands for the Lieb lattice on a path between the symmetry points of the Brillouin zone, see Tab. 1 for the parameter values used.
@@ III Electron model
where $c_{i\sigma}^{({\dagger})}$ annihilates (creates) an electron with spin $\sigma$ at site $i$. The pairs $\langle i,j\rangle$ and $\langle\langle i,j\rangle\rangle$ denote nearest- and next-nearest neighbors with hopping amplitudes $t$ and $t_{2}$, respectively. $J_{\mathrm{sd}}$ is the on-site exchange coupling between the itinerant electrons and the localized spins $\textbf{S}_{i}$, and $\bm{\sigma}$ is the vector consisting of Pauli matrices. The spins $\textbf{S}_{i}$ have a length $S$, and the on-site energy is set by the chemical potential $\mu$. We assume that spin-orbit coupling is negligible compared to the spin-splitting.
The electronic energy spectrum obtained from this Hamiltonian is showcased in Fig. 2 (a), which exhibits the characteristic $d_{x^{2}-y^{2}}$-wave spin-split electron bands of an altermagnet. The key feature when finding the spin-polarized superconductivity is the existence of regions that are completely spin-split for all momenta, which can be observed between the middle and top bands. To obtain this spectrum, we diagonalize the Hamiltonian in momentum space as $H_{\mathrm{el}}=\sum_{n\textbf{k}\sigma}\varepsilon_{n\textbf{k}\sigma}d^{\dagger}_{n\textbf{k}\sigma}d_{n\textbf{k}\sigma}$, where $\varepsilon_{n\textbf{k}\sigma}$ is the energy of band $n$ for excitations with momentum k and spin $\sigma$, while $d^{\left({\dagger}\right)}_{n\textbf{k}\sigma}$ are the annihilation (creation) band operators. These operators are related to the electron operators as $d_{n\textbf{k}\sigma}=\sum_{\alpha}q_{\textbf{k}\sigma\alpha n}^{*}c_{\textbf{k}\sigma\alpha}$, where $\alpha$ indicates the different atoms in the basis, and the coefficient $q_{\textbf{k}\sigma\alpha n}$ is the component $\alpha$ in the eigenvector of the Hamiltonian corresponding to band $n$.
Figure 2: (a) The electronic band structure along a high-symmetry path in the BZ for $\mu=0$. (b) The FSs for spin-up (red) and spin-down (blue), the colors indicate $at\left|d\varepsilon/dk_{\perp}\right|^{-1}$ for $\mu=0.1t$. The NNN hopping is $t_{2}\approx 0.21t$ for spin-up, and $t_{2}=0$ for spin-down. This particular system exhibits altermagnetism with $d_{x^{2}-y^{2}}$-wave spin splitting. See Tab. 1 for the rest of the parameters used.
Due to the $d_{x^{2}-y^{2}}$-wave symmetry of the spin splitting, spin-up and spin-down bands are interchangeable by a rotation of $90^{\circ}$ in momentum space. Hence, all properties of the two spins are connected by this rotation. To showcase the role of NNN hopping, we present the results with a NNN hopping of $t_{2}\approx 0.21t$ for spin-up electrons and the results with $t_{2}=0$ for spin-down electrons throughout the paper. The electronic band structure for $t_{2}=0$ can be found in Ref. [39].
@@ IV Electron-phonon coupling
where $\textbf{r}_{i\alpha j\beta}$ is the instantaneous separation between atoms $i\alpha$ and $j\beta$, $\bm{\delta}$ are vectors connecting the atoms in equilibrium, and $\textbf{u}_{ij\alpha\beta}$ is their relative displacement from equilibrium. $\bm{\nabla}_{r}$ denotes the gradient with respect to $\textbf{r}_{i\alpha j\beta}$, and $|_{\textbf{r}_{ij\alpha\beta}=\bm{\delta}}$ indicates to evaluate the atom positions in equilibrium. See App. C for the full derivation of the electron-phonon coupling. The first term of (3) is the equilibrium hopping used in (2), while the second term is the source of the electron-phonon interaction. We quantize the displacements as [75, 76]
$\textbf{u}_{i\alpha}=\sum_{\textbf{q}\lambda}\sqrt{\frac{\hbar}{2M_{\alpha}N\omega_{\textbf{q}\lambda}}}\bm{\hat{e}}_{\textbf{q}\alpha}^{\lambda}\left(a_{-\textbf{q},\lambda}^{\dagger}+a_{\textbf{q}\lambda}\right)e^{-i\textbf{q}\cdot\textbf{R}_{i\alpha}},$ (4)
where $a_{\textbf{q}\lambda}^{({\dagger})}$ annihilates (creates) a phonon with momentum q in mode $\lambda$, $N$ is the number of unit cells, and $\textbf{R}_{i\alpha}$ is the equilibrium position of atom $\alpha$ in unit cell $i$.
@@
$\times\sum_{\bm{\delta}_{j}}e^{i\textbf{k}\cdot\bm{\delta}_{j}}\left(\frac{\delta_{jx}}{\sigma_{x}^{2}},\frac{\delta_{jy}}{\sigma_{y}^{2}}\right)\cdot\left[\frac{\bm{\hat{e}}_{\textbf{q}\beta}^{\lambda}}{\sqrt{M_{\beta}}}e^{i\textbf{q}\cdot\bm{\delta}_{j}}-\frac{\bm{\hat{e}}_{\textbf{q}\alpha}^{\lambda}}{\sqrt{M_{\alpha}}}\right],$
where $j\in\{1,2\}$ indicates whether it is the NN or the NNN coupling, and $\sigma_{x},\sigma_{y}$ are the standard deviations of the atomic orbitals. Note that we have omitted the band indices, as we consider the scattering to be confined to the FS.
@@ V Superconductivity
Next, we transform the electron-phonon interaction into an effective electron-electron interaction by applying a Schrieffer-Wolff transformation [80, 81], see App. C for the full derivation,
$H_{\mathrm{eff}}=\frac{1}{2}\sum_{\textbf{k}\textbf{k}^{\prime}\sigma}\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{\textbf{k}^{\prime}\textbf{k}\sigma}d^{\dagger}_{\textbf{k}^{\prime}\sigma}d^{\dagger}_{-\textbf{k}^{\prime},\sigma}d_{-\textbf{k},\sigma}d_{\textbf{k}\sigma},$ (7)
where we have followed Refs. [34, 58], and symmetrized the effective interaction as $2\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{\textbf{k}^{\prime}\textbf{k}\sigma}\equiv V_{\textbf{k}^{\prime}\textbf{k}\sigma}+V_{-\textbf{k}^{\prime},-\textbf{k}\sigma}-V_{-\textbf{k}^{\prime},\textbf{k}\sigma}-V_{\textbf{k}^{\prime},-\textbf{k}\sigma}$, with the effective electron-electron interaction on the FS of spin $\sigma$, $V_{\textbf{k}^{\prime}\textbf{k}\sigma}$ given by
@@
As shown in Refs. [39, 58] the electrons have the greatest DOS for $J_{\text{sd}}S<\mu<J_{\text{sd}}S$ so we expect the greatest $T_{c}$ here. In the region $0<\mu<J_{\text{sd}}S$, there is no overlap of the spin up and spin down FSs, also with $t_{2}\neq 0$. Hence, spin-singlet zero-momentum pairing is ruled out, and we focus on this region. The most likely scenario with the type of FS of our model appears to be zero-momentum pairing. This provides the greatest possible phase space for interactions, and, unlike other models, there is a very small overlap of the spin up and spin down FSs even with a finite momentum, App. E. Hence, we look for spin-triplet pairing, which necessitates odd-parity gap functions, and hence an odd-parity symmetrized coupling [16, 34]. The symmetrized interaction is shown in Fig. 3 (a) and depicts a $p$-wave symmetry.
We consider weak-coupling superconductivity and apply the linearized BCS equation to solve for the spin-polarized superconducting gap $\Delta_{\textbf{k}\sigma}$ and critical temperature. Since electrons close to the FS dominate the superconducting pairing, we utilize the FS average of the BCS equation [39, 58, 82, 83]